Part D Without actually calculating the logarithm, determine what two integers the value of log 1.37x O9 falls between. The value of log(1.37 x 10) falls between Please Choose Please Choose 0 and 1 1 and 2 7 and 8 8 and 9 9 and 10 Submit Request Answer Part E Complete previous partís)
Answer #1

Remember these logarithm properties for these type of questions:

log (m*n) = log m + log n

log (m/n) = log m - log n

log (m)n = n*log m

Value of log 10 = 1

Value of log 1 = 0

So Now given that

log (1.37*109)

Using 1st property (where m = 1.37 & n = 109)

= (log 1.37) + log 109

Now Using 3rd property

= (log 1.37) + 9*log 10

(Since log 10 = 1, So)

= log (1.37) + 9*1

= 9 + log (1.37)

(since log 1 = 0, & log 10 = 1, In between these (for 1.37) value of log will be less than 1)

= 9 + something less than 1 and greater than 0

So the value of log (1.37*109) falls between 9 and 10.

Please Upvote.

Comment below if you have any query.